OpenStudy (anonymous):

A new ATV worth $500 is depreciating in value by $50 per year. The mathmatical model y=-50x+500 describes the value, y, in dollars after x years. Find the ordered pairs which represents the value of the ATV after 8 years.

OpenStudy (anonymous):

you have to put the number after 8 years for example you have to put 8 in model y=-50*8+500=100 then you have to put 9 in model y=-50*9+500= 50 then you have to put 10 in model y=-50*10+500=0 So the price is getting to zero So we have 3 order pairs of (8,100) and (9,50) and (10,0)

OpenStudy (anonymous):

i think you have to go one for other years as well